OpenStudy (anonymous):

what percent of 25 is 23

OpenStudy (eyust707):

use the same method

OpenStudy (anonymous):

92% On a calculator, divide 23 by 25 and then multiply by 100 a percent is 1/100 23/25 = 92/100 = 92%

OpenStudy (eyust707):

(23/25) *100

OpenStudy (anonymous):

hope it know

OpenStudy (anonymous):

woops i meant i hope you get it! :)

OpenStudy (anonymous):

Thanks but im not sure what to type in because all of these problems are differnt

OpenStudy (anonymous):

different*